# Polish Academy of Sciences

### v1

@@ Abstract
The article presents an innovative approach to the visualisation of multidimensional data, using icons inspired by Chernoff faces. The approach merges classical projection techniques with the assignment of particular data dimensions to mimic features, capitalizing on the natural ability of the human brain to interpret facial expressions. The technique is implemented as a plugin to the dpVision open-source image handling platform. The plugin allows the data to be interactively explored in the form of a swarm of ”totems” whose position in hyperspace as well as facial features represent various aspects of the data. Sample visualisations, based on synthetic test data as well as the vinhoverde 15-dimensional database on Portuguese wines, confirm the usefulness of our approach to the analysis of complex data structures.
Keywords: Multidimensional data, data visualisation, Chernoff faces, avatar glyphs, interactive 3D, dimensionality reduction, PCA, human-centred visualisation, dpVision
@@ 1 Introduction
    increase the dimensionality of the display by bestowing the representations of the points with additional characteristics besides their coordinates. Those additional properties can include color, size, shape parameters, any quantifiable pecularities of appearance. Each data point is then represented by a more elaborate token, which in this text will be referred to as an avatar. One well-known approach is the use of Chernoff faces [1], icons resembling a human face, with data values converted into parameters such as rounded/oblong shape of the face, hair color and length, smiling/frowning curvature of lips etc. One advantage of Chernoff faces is that they tap into the natural ability of the human observer to identify and interpret human facial traits and expressions. However, this same capacity introduces challenges, as human observers may be influenced by the emotional connotations carried by various types and expressions of human or human-like faces. Perceiving an avatar as attractive/ugly or hostile/friendly adds a factor which is not necessarily related to the actual data and can seriously distort the analysis. To overcome this problem, Lardelli [2] proposed alternative glyphs in the form of trees. Trees are very familiar to humans, but seldom evoke strong emotions.
Both approaches will be used in this work.
The type of multivariate data taken into consideration will be a cloud (finite set) of points in a multidimensional hyperspace.
The multidimensional datasets chosen for testing purposes are a) an artificial 4D shape chosen for easy visual interpretation, and b) the publicly available Vinhoverde set of data on Portuguese wines [3].
2 Symbols table
$K$
number of data points;
$n$
number of dimensions in the data to be visualized;
$m<n$
the number of dimensions the avatars are capable of representing by appearance;
$h\leq n$
number of data dimensions meaningful to the human user. By this we mean parameters which the user can relate to while thinking about the data set, because they are relevant to her work or her perception of datapoints as advantageous/disadvantageous:
$h_{s}\leq h$
number of those among them that were chosen for presentation as hyperspacial dimensions;
$h_{v}=h-h_{s}$
number of those among them that were chosen for presentation by avatar appearance;
$k\geq 3$
number of display dimensions the user is prepared to navigate. In our experiments, we limited this number to 4, so as not to overwhelm the user.
@@ 3 State of the art
The parallel coordinate plot [4] reduces the n-dimensional coordinate system to the 2D display plane by placing the coordinate axes in parallel (usually in a vertical direction) and representing each nD data point as a polyline with individual coordinate values as its nodes (Fig. 2 (b, c)). An early application of this format in medicine was known as liver profiles, representing a single nD (typically, 6D) point as a polyline or as a set of parallel vertical bars of various height (Fig. 2 (a)). When multiple nD points need to be displayed, the polylines are colored to keep them distinguishable (Fig. 2 (b)). A smoothed version of this construct is the Andrews plot (Fig. 2 (c)) [5].
A similar approach in a different (radial) coordinate system uses radial coordinates (Fig. 2 (d)). While carrying the same information as parallel coordinates, its advantage is a somewhat easier visual comparison of plots, especially in relatively low dimensionalities (up to about 10).
@@
A trellis chart (Fig. 3) is a compact array of small 2D plots (usually all of the same type), showing relationships between two dimensions at a time, allowing the user or reader to examine them in a more comprehensive way than if they were spread over several pages or screens. A classic setup is a full set of all (${N\choose{2}}=\frac{n(n-1)}{2}$ possible pairs of dimensions, fitted into one page or screen.
@@
Purdom [6] provides a review of multivariate presentation, going through the four types of variables: nominal (categorical), ordinal, discrete and continuous. Her paper includes large format trellis charts, high-resolution maps using color for a third dimension, and the use of Principal Component Analysis (PCA) to choose linear combinations of variables as aggregate ”dimensions” representing the most pertinent aspect of the data.
@@
A more condensed representation was introduced by B. Alpern and L. Carter [7]. They represent the axes of an n-dimensional coordinate system as a 2D pencil of line segments emanating from a common origin and contained in an angle of less than 180 degrees. $n-1$ additional copies of each segment are used to form a mesh of ($N\choose{2}$ paralellograms, representing the coordinate planes of the hyperspace in a hyperbox (Fig. 4), displayed with the physical appearance of a mutifaceted 3D convex polyhedron in perspective.
Despite the large number of methods proposed for the visualisation of multidimensional data, most of them focus either on mathematical reduction of dimensionality (such as PCA or t-SNE), or on geometric representations like parallel coordinates, radial plots, or static glyphs. These methods typically treat all data dimensions equally, without considering how understandable each dimension is to the human observer.
In contrast, our approach introduces two important elements that, to our knowledge, are not present in previous work. First, we clearly separate the dimensions of the dataset into two groups: those that are intuitively meaningful to users (such as quality ratings or sensory features), and those that are more abstract or technical (such as chemical coefficients). The intuitive features are assigned to visual characteristics of avatars — inspired by Chernoff faces — while the abstract features are reduced or projected into 3D spatial coordinates.
Second, we combine this concept with interactive 3D visualisation, allowing users to explore a swarm of avatars from different angles and positions. While face-based visualisation is known from Chernoff’s work, it has mostly been used in two-dimensional form. We are not aware of any previous system that integrates avatar-based glyphs with interactive spatial navigation and dimensional filtering in a 3D scene.
@@ 4 Paradigm
Quite often, when multidimensional data is mentioned, we think of abstract sets of numbers to be analyzed by pure mathematics. However, there do exist data sets in which at least a part of the parameters is intuitively recognizable and can be directly interpreted in the context of a given subject area. One example is the publicly available dataset describing Portuguese wines, where find such parameters as acidulity, sugar content, or general assessment of quality. These features are readily understandable both to the expert and to the consumer. We can therefore divide multidimensional data into two categories.
The first group is those parameters that are intuitively understandable to the human observer. The values of parameters in this group refer to perceptible qualities of the objects in question, which makes them easier to interpret and compare. In the case of wines it’s the sensory and quality attributes mentioned above, to which the user can relate from his own experience. Displaying them individually as separate dimensions provides for a better analysis and understanding of the data structure.
The other group includes those parameters, which, however important they can be mathematically, have no direct reference to human perception. They can be chemical coefficients or technical aspects of the manufacturing process, whose meaning is difficult to perceive without additional context. In such cases the computer can be used to reduce their dimensionality to a smaller number of components, preserving the most essential information while reducing the burden on the visual channel.
This conceptual division between user-recognisable and abstract dimensions is central to our method. In most visualisation systems, all dimensions are treated as equal data fields, often without considering how understandable they are to the user. While this is sufficient in purely mathematical applications, it may create problems in real-world contexts, where users expect visualisations to reflect qualities they can interpret and relate to.
By assigning the more intuitive dimensions to the appearance of avatars, we use the natural strength of human perception — the ability to notice and compare visual traits. The less intuitive dimensions are projected into 3D space using classical techniques or reduced using PCA, allowing the user to explore general patterns and clusters without being distracted by unfamiliar technical details.
This dual approach makes the system flexible and adaptable to different data types and user needs. It also reflects a real-world situation, where some features (like sweetness or colour of a wine) are easier to explain and compare, while others (like pH level or sulphate concentration) may require expert interpretation. Our system helps to balance these aspects in a single coherent visualisation.
This approach merges the advantages of two tools of analysis: on one hand, the user is offered a clear and intuitive representation of the most meaningful data fields; on the other — the less intuitive ones are filtered by the computer, finally yielding a visualisation which is both information-rich and easy to interpret.
And this representation will be delivered to the user via two faculties of the human visual system: our capability of assessing spatial relationships in a 3D space, and our aptitude for perceiving facial expressions, although in a less quantitative way.
The data dimensions which were chosen for spatial display are transformed by two matrices and one scalar parameter:
  • $\mathbf{F}$
    filtering matrix which incorporates the following functions:
    1. 1.
      Selection of named parameters;
    2. 2.
      If necessary, transformation of unnamed parameters into a reduced space (PCA);
    3. 3.
      assignment of both groups to dimensions of observation space;
  • $\mathbf{V}$
    viewing matrix, describing the position and optics of the observer moving around in observation space. This matrix and the accompanying viewing algorithm determine which data points should actually be visible in the 3D viewing space, and at what position.
  • $s$
    slab thickness is the tolerance to which a data point must intercept the viewer hyperplane to be displayed. It can be set to infinity, in which case all data points will be projected onto the hyperplane.
@@
The visualisation procedure takes as its input:
  1. 1.
    the number $n$ of dimensions of the cloud;
  2. 2.
    an $n*K$ data cloud (a matrix in which each of k data points is a column, and each dimension is a row);
  3. 3.
    a transformation matrix with $n$ columns and $m+k$ rows, $m$ being the number of visual parameters of an avatar.
  4. 4.
    the $(k+2)*n$ observer matrix. It is used to transform data from the $n$-dimensional data space into the k-dimensional observation space.
    Figure 5: Fragment of dialog box assigning parameters
    These transformations, projecting the data points both into the geometric visualization hyperspace and into the parameter space of the Chernoff faces, are computed from the user’s choices made in a dialog box (Fig. 5), assigning various display forms to data dimensions. Some data dimensions can be left unassigned, and they will be subjected to PCA analysis, to reduce their number to that of the available spacial dimensions.
    One of the two additional rows, the slice thickness is used to determine the distance between a given data point and the crossection hyperplane (points too far away from it are not visualized). The other extra row is needed for homogeneous representation of the data.
  5. 5.
    the viewing matrix, part of which is controlled by the user via mouse or keyboard commands, using a multidimensional compass rose to visualize the current orientation and shift. Another user-controlled parameter is the thickness of the visible slab (area between two parallel hyperplanes).
Figure 6: A point cloud intersected by a slab
Upon completion, the procedure inserts a cloud of avatars into the viewing space, i.e. the working space of dpVision [8, 9] (see Fig. 7).

### v2

@@ Abstract
The article presents an innovative approach to the visualization of multidimensional data, using icons inspired by Chernoff faces. The approach merges classical projection techniques with the assignment of particular data dimensions to mimic features, capitalizing on the natural ability of the human brain to interpret facial expressions. We introduce a semantic division of data dimensions into intuitive and technical categories, assigning the former to avatar features and projecting the latter into a hyperspace of four, or potentially more dimensions. The technique is implemented as a plugin to the dpVision open-source image handling platform. The plugin allows the data to be interactively explored in the form of a swarm of avatars whose position in hyperspace as well as facial features represent various aspects of the data. Sample visualizations, based on synthetic test data as well as the 12-dimensional database on Portuguese Vinho Verde wines, confirm the usefulness of our approach to the analysis of complex data structures.
@@ 1 Introduction
    increase the dimensionality of the display by bestowing the representations of the points with additional characteristics besides their coordinates. Those additional properties can include color, size, shape parameters, any quantifiable visual peculiarities. Each data point is then represented by a more elaborate token, which in this text will be referred to as an avatar. One well-known approach is the use of Chernoff faces [1], icons resembling a human face, with data values converted into parameters such as rounded/oblong shape of the face, hair color and length, smiling/frowning curvature of lips etc. One advantage of Chernoff faces is that they tap into the natural ability of the human observer to identify and interpret human facial traits and expressions. However, this capacity also introduces challenges, as human observers may be influenced by the emotional connotations carried by various types and expressions of human or human-like faces. Perceiving an avatar as attractive/ugly or hostile/friendly adds a factor which is not necessarily related to the actual data and can seriously distort the analysis. To overcome this problem, Lardelli [2] proposed alternative glyphs in the form of trees. Trees are very familiar to humans, but seldom evoke strong emotions.
In this work, we propose a novel approach that semantically separates data dimensions into two groups: those that are intuitively meaningful to the user (e.g., sensory, categorical, interpretable parameters), and those that are abstract or technical. Users may assign meaningful dimensions to either geometric space or avatar visual features. If fewer than the required number of dimensions are assigned, the system automatically supplements them using anonymous dimensions, which are processed via PCA to generate missing spatial or visual features. This approach is implemented as an interactive plugin for the dpVision platform [3, 4], allowing users to explore a cloud of 3D avatars positioned in a 4D data space and visualized through dynamic projections into 3D space (see Figure 1). By merging spatial navigation with avatar-based representation, our plugin provides a human-centered and intuitive visualization of complex multidimensional data.
@@
To validate the proposed approach, we use two types of multivariate datasets: a) an artificial 4D structure designed for easy geometric interpretation, and b) the publicly available dataset containing 12-dimensional data on Portuguese Vinho Verde wines [5, 6].
@@ 2 STATE OF THE ART
The parallel coordinate plot [7] reduces the n-dimensional coordinate system to the 2D display plane by placing the coordinate axes in parallel (usually in a vertical direction) and representing each nD data point as a polyline with individual coordinate values as its nodes (Figure 2 (b, c)). An early application of this format in medicine was known as liver profiles, representing a single nD (typically, 6D) point as a polyline or as a set of parallel vertical bars of various height (Figure 2 (a)). When multiple nD points need to be displayed, the polylines are colored to keep them distinguishable (Figure 2 (b)). A smoothed version of this construct is the Andrews plot (Figure 2 (c)) [8].
A similar approach in a different (radial) coordinate system uses radial coordinates (Figure 2 (d)). While carrying the same information as parallel coordinates, its advantage is a somewhat easier visual comparison of plots, especially in relatively low dimensionalities (up to about 10).
@@
A trellis chart (Figure 3) is a compact array of small 2D plots (usually all of the same type), showing relationships between two dimensions at a time, allowing the user or reader to examine them in a more comprehensive way than if they were spread over several pages or screens. A classic setup is a full set of all $\binom{N}{2}=\frac{n(n-1)}{2}$ possible pairs of dimensions, fitted into one page or screen.
@@
Purdom [9] provides a review of multivariate presentation, going through the four types of variables: nominal (categorical), ordinal, discrete, and continuous. Her paper includes large format trellis charts, high-resolution maps using color for a third dimension, and the use of Principal Component Analysis (PCA) to choose linear combinations of variables as aggregate "dimensions" representing the most pertinent aspect of the data.
@@
A more condensed representation was introduced by B. Alpern and L. Carter [10]. They represent the axes of an n-dimensional coordinate system as a 2D pencil of line segments emanating from a common origin and contained in an angle of less than 180 degrees. The $n-1$ additional copies of each segment are used to form a mesh of $\binom{N}{2}$ parallelograms, representing the coordinate planes of the hyperspace in a hyperbox (Figure 4), displayed with the physical appearance of a multifaceted 3D convex polyhedron in perspective.
Despite the large number of methods proposed for the visualization of multidimensional data, most of them focus either on mathematical reduction of dimensionality (such as PCA or t-SNE), or on geometric representations like parallel coordinates, radial plots, or static glyphs. These methods typically treat all data dimensions equally, without considering how understandable each dimension is to the human observer.
An alternative, cognitively grounded approach was proposed by Pflughoeft et al. [11], who introduced data avatars — metaphorical figures designed to visually encode multidimensional data in ways that align with human intuitive perception. Their work, based on dual-process cognitive theory, shows that representations tailored to intuitive understanding can improve user performance in data interpretation tasks.
We propose an approach that introduces two complementary elements not present in previous work. First, we clearly separate the dimensions of the dataset into two groups: those that are intuitively meaningful to users (such as quality ratings or sensory features), and those that are more abstract or technical (such as chemical coefficients). The intuitive features are assigned to visual characteristics of avatars—inspired by Chernoff faces—while the abstract features are reduced or projected into 3D spatial coordinates.
Second, we combine this concept with interactive 3D visualization, allowing users to explore a swarm of avatars from different angles and positions. While face-based visualization is known from Chernoff’s work, it has mostly been used in two-dimensional form. We are not aware of any previous system that integrates avatar-based glyphs with interactive spatial navigation and dimensional filtering in a 3D scene.
3 NOTATION AND TERMINOLOGY
Throughout this paper, we use consistent notation to describe data dimensions, transformation matrices, and visualization parameters. The key symbols and abbreviations are summarized in Table 1, and are referenced throughout the description of the projection procedure and implementation details.
Table 1: Symbols and notation used throughout the paper.
Symbol Meaning Constraints
Input data parameters
$N$ Total number of data points
$n$ Total number of data dimensions
$X$ Original data matrix $X\in\mathbb{R}^{n\times N}$
Implementation-dependent mapping
$k$ Number of navigable display dimensions $k\geq 3$ (typically 4)
$m$ Number of available avatar visual features $m\leq n-k$ (currently 10)
User’s mapping
$h_{S}$ Number of human-meaningful dimensions mapped to space $h_{S}\leq k$
$h_{V}$ Number of human-meaningful dimensions mapped to avatars $h_{V}\leq m$
$h_{A}$ Number of anonymous dimensions $h_{A}\leq n-(h_{S}+h_{V})$
Projection matrices
$F$ Filtering matrix $F\in\mathbb{R}^{(k+m)\times n}$
$V$ View matrix $V\in\mathbb{R}^{(k+1)\times(k+1)}$
Derived datasets
$X_{\text{spatial}}$ Spatial projection of $X$ $X_{\text{spatial}}\in\mathbb{R}^{k\times N}$
$X_{\text{visual}}$ Avatar feature mapping of $X$ $X_{\text{visual}}\in\mathbb{R}^{m\times N}$
$X_{\text{view}}$ Spatial data after applying $V$ $X_{\text{view}}\in\mathbb{R}^{k\times N}$
@@ 4 PARADIGM
Multidimensional data sets often contain a mixture of parameters: some of them intuitively understandable to humans, and others more abstract or technical. To maximize interpretability, we propose a semantic division of data dimensions into two groups:
  • Intuitive dimensions, such as quality ratings or sensory properties, are mapped onto visual characteristics of avatars (e.g., face shape, expressions, color).
  • Technical dimensions, such as chemical or structural parameters, are projected into spatial coordinates in a 3D observation space after optional dimensionality reduction.
Figure 5: Conceptual diagram illustrating the separation of multidimensional data into spatial coordinates and avatar visual features through filtering and viewing transformations.
Let the original data set be represented by a matrix $X\in\mathbb{R}^{n\times N}$, where $n$ is the number of original dimensions and $N$ is the number of data points.
A filtering matrix $F\in\mathbb{R}^{(k+m)\times n}$ is constructed to separate and process the dimensions, where $k$ is the number of spatial dimensions and $m$ the number of visual avatar features. The filtered data are:
$X_{\text{filtered}}=F\cdot X$
Here, each row of $F$ defines a new output feature (spatial or visual) as a linear combination of the original dimensions of $X$.
Spatial coordinates are obtained from a subset $F_{\text{spatial}}$, where $F_{\text{spatial}}\in\mathbb{R}^{k\times n}$, resulting in:
$X_{\text{spatial}}=F_{\text{spatial}}\cdot X$
Similarly, visual parameters are derived from $F_{\text{visual}}$, where $F_{\text{visual}}\in\mathbb{R}^{m\times n}$
$X_{\text{visual}}=F_{\text{visual}}\cdot X$
The spatial data are further transformed using a viewing matrix $V\in\mathbb{R}^{k\times n}$ (where typically $k=3$ or $4$):
$X_{\text{view}}=V\cdot X_{\text{spatial}}$
This dual mapping—avatars for intuitive features, spatial positioning for abstract features—provides a flexible and human-centered way to visualize high-dimensional data.
The conceptual framework of the proposed method, including the separation of intuitive and technical dimensions and the application of the filtering and viewing matrices, is illustrated in Figure 5.
@@ 5 IMPLEMENTATION
Based on the conceptual framework presented in the previous section, the practical implementation of the proposed visualization method involves the construction of appropriate transformation matrices, dimensionality reduction where necessary, and the assignment of selected data dimensions to spatial coordinates or avatar visual features.
The implementation uses the Eigen library (C++), which allows efficient operations on matrices and vectors.
The input data are stored in a matrix of size $n\times N$:
Eigen::MatrixXd X(n, N);
Each column corresponds to a single point in the data space $\mathbb{R}^{n}$, and each row represents one of the $n$ dimensions in that space.
5.1 Assigning dimensions to categories
Note: To avoid confusion, we refer to dimensions when describing input data, and to features when referring to the visualized output.
Figure 6: Dialog window for assigning input dimensions to feature categories: spatial, visual, unnamed, or skipped. For spatial and visual categories, a specific output feature label can be selected.
Input dimensions can be assigned to one of three feature categories: spatial, visual, or anonymous. Additionally, any dimension may be marked as skipped and excluded from further processing.
This assignment is performed through a dedicated dialog window (Figure 6), which enables the user to associate input dimensions with specific feature categories and, when applicable, select predefined visual or spatial feature labels.
The complete set of available labels for spatial and visual features is presented in Table 2.
The resulting configuration is stored in index tables. It is not required to assign all available features — any missing spatial or visual values are automatically interpolated during subsequent processing.
Table 2: List of features used in the implementation
Category Feature Description
User-defined spatial features
spatial X X coordinate in 4D space
spatial Y Y coordinate in 4D space
spatial Z Z coordinate in 4D space
spatial T T coordinate in 4D space
User-defined avatar facial attributes
visual Skin_C avatar’s skin color
visual Hair_C avatar’s hair color
visual Eye_S distance between the avatar’s eyes
visual Nose_L avatar’s nose length
visual Mouth_W avatar’s mouth width
visual Smile degree of curvature of the avatar’s smile
visual Frown level of frowning expression
visual Hair_L avatar’s hair length
visual Face_Elong avatar’s face proportions
visual Iris_C avatar’s iris color
features not defined by the user
anonymous used for principal component analysis (PCA)
Features selected by the user to skip
skipped excluded from further processing
5.2 Filtering matrix
The first step is to create the filtering matrix, which in the implementation has the following form:
Eigen::MatrixXd F(k + m, n);
F.setZero();
To simplify later operations, we assume that the first $k$ rows of this matrix correspond to spatial features (in our implementation, these are the 4 features: X, Y, Z, T), and the following $m$ rows correspond to visual features (in our case: 10 features representing details of the avatar’s appearance). The matrix $F$ is initially filled with zeros and then completed based on the feature assignments:
  • First, for each of the $h_{S}$ spatial features and $h_{V}$ visual features selected by the user, 1 is set at the intersection of the row for that feature and the column of the original input dimension.
  • Then, if there are still unused rows in $F$, and the user has selected anonymous features ($h_{A}>0$), a PCA is performed on the input dimensions assigned to those anonymous features. The $(k+m)-(h_{S}+h_{V})$ strongest components are selected, and their coefficient vectors are inserted into the remaining empty rows of $F$.
The transformed data is calculated as the product of the filtering matrix and the input matrix, followed by normalization. The procedure for creating the filtering matrix and transforming the input data is shown in Algorithm 1.
Algorithm 1 Implementation of Data Projection
0:  $X$ (all features), $Asgn$ (assignment of features to groups)
0:  $X_{\text{filtered}}$ (projected data)
1:  Initialize $F$ (zero matrix)
2:  for each feature index $j$ do
3:     if assignment of feature $j$ is ignored then
4:        Skip
5:     else if assignment of feature $j$ is spatial then
6:        Mark corresponding spatial row in $F$
7:     else if assignment of feature $j$ is visual then
8:        Mark corresponding visual row in $F$
9:     else if assignment of feature $j$ is anonymous then
10:        Save index of feature $j$ for PCA
11:     end if
12:  end for
13:  if number of mapped features $(h_{S}+h_{V})<(k+m)$ then
14:     Perform SVD on anonymous features
15:     Select $(k+m)-(h_{S}+h_{V})$ top components
16:     Insert principal directions into empty rows in $F$
17:  end if
18:  $X_{\text{filtered}}=F\times X$
19:  for each row in $X_{\text{filtered}}$ do
20:     Center and scale (standardize)
21:     if row corresponds to spatial dimension then
22:        Shift by $+0.5$
23:     end if
24:  end for
25:  return  $X_{\text{filtered}}$
In this way, we obtain a table of $N$ points, each with $(k+m)$ dimensions. These dimensions can be divided into two categories:
  • The first $k$ rows form the matrix X_spatial, which contains spatial features.
  • The following $m$ rows form the matrix X_visual, which contains avatar visual features.
Eigen::MatrixXd X_filtered(k + m, N);
X_filtered = F * X;
Eigen::MatrixXd X_spatial = X_filtered.topRows(k);
Eigen::MatrixXd X_visual = X_filtered.bottomRows(m);
The matrix $\mathbf{X}_{\text{spatial}}$ is now a ready-to-use $k$-dimensional dataset for further processing.
5.3 View matrix
The implementation supports navigation in a $k$-dimensional space (typically 4D) using affine transformations. These transformations are represented by a square matrix of size $(k+1)\times(k+1)$, which combines both rotation and translation.
The core of this matrix is a $k\times k$ rotation matrix, generated by specifying the rotation plane (two coordinate axes) and the rotation angle.
In the current plugin implementation, all 4D rotations are carried out in coordinate planes such as XY, XZ, XT, YZ, YT, or ZT. These correspond to classic SO(2) rotations within two-dimensional subspaces. However, the implementation can be extended to support full SO(4) rotations, offering a more expressive and flexible navigation model.
The remaining part of the view matrix stores a translation vector and a homogeneous row to enable affine transformations.
In code, the matrix can be built as follows:
// Rotation only (k x k)
Eigen::MatrixXd R = generateRotationMatrix(k, axis1, axis2, angle);
// Create full (k+1 x k+1) view matrix with translation
Eigen::MatrixXd V = Eigen::MatrixXd::Identity(k + 1, k + 1);
V.topLeftCorner(k, k) = R;
V.topRightCorner(k, 1) = translation_vector; // optional
To apply the transformation, the spatial data must be converted into homogeneous coordinates by adding a row of ones:
Eigen::MatrixXd X_hom(k + 1, N);
X_hom.topRows(k) = X_spatial;
X_hom.row(k) = Eigen::RowVectorXd::Ones(N);
Eigen::MatrixXd X_view = V * X_hom;
This allows smooth 4D navigation, including both rotation and panning of the data in the observation space.
5.4 Slab-based Filtering
To limit the number of visible data, a slab-based filtering mechanism is used. The goal is to retain only those data points that lie close to the viewing space—within a given distance slab_thickness from a hyperplane defined by the current rotation.
The filtering uses a normal vector to the viewing space to compute signed distances from each point. Depending on whether the filtering is applied before or after transforming the data by the view matrix, different normal vectors are used:
  • if filtering is done in the original spatial coordinates (X_spatial), the normal vector can be taken as the last row of the rotation matrix, which defines the orientation of the view;
  • if filtering is done after applying the rotation (X_view), the normal vector can be defined in canonical coordinates as $[0,0,0,1]$, which corresponds to the last axis.
Each point is projected onto the normal vector, and only those within the allowed threshold are retained.
Eigen::VectorXd slab_values = normal_vector.transpose() * X_view;
auto mask = slab_values.cwiseAbs().array() < slab_threshold;
Eigen::MatrixXd X_visible = skip_masked(X_view, mask);
Points that satisfy the slab condition—meaning they are sufficiently close to the viewing hyperplane—are passed to the next stage of the pipeline. This filtering effectively creates a thick slice through the multidimensional space (Figure 7).
Figure 7: A cloud of points intersected by a slab, using a threshold of 1.5. Red points that lie outside the slab are discarded.
5.5 3D projection
The four-dimensional coordinates are then projected into 3D space using a standard perspective projection. Avatars are generated as three-dimensional objects based on visual features stored in the $\mathbf{X}_{\text{visual}}$ matrix and rendered using the tools provided by the dpVision platform (Figure 8). To ensure interactive performance with large datasets, the avatars are deliberately simplified in geometry, allowing a large number of them to be rendered and manipulated in real time.